LFBG Re-Commissioning with LUCI: 21 September 2010 UT

Observer: DThompson, ARakich, DMiller (Tucson), JHill (LBTO)
Telescope Operator: SAllanson and GBechetti (LBTO)
Telescope Support: JLittle, MMidkiff (LBTO)
Instrument Support: DOfficer (LBTO)

Summary

Some monsoon-like showers threatened early, but they stayed to the South, so we worked all night with the occasional low clouds drifting by. We worked all night on field aberration measurements to adjust the pointing of M2 and M3 to minimize binodal astigmatism and focal plane tilt. Andrew and Doug converged on a good alignment for M2 and M3 which we will now make permanent, and they took some new calibration data for both M2 and M3. This completes the last critical measurement for LFBG re-commissioning. Dave made the first ever range-balancing operation by Mode 1 repointing M1 and M2 in morning twilight. Seeing was 0.4 to 0.6 arcsec for most of the night. Had to reset the LFBG rotator, but the telescope was otherwise well-behaved.

Adopted the following Left Arbitrator Pointing Settings which will be absorbed into the collimation models:
M2 Tip=+9 Tilt=-35 arcsec which correspond to IE=-81 CA=+20
M3 Tip=+75 Tilt=0 arcsec (absolute -540 Tip -20 Tilt)
Note: These adjustments have been absorbed into the global offsets for LUCI and IRTC the following evening.

Preparations

Details

Open

01:52 Open to partly cloudy sky

Pointing Check

*** Pointing Model Terms ***
IA      -406.41010
IE       -24.92430
NPAE       0.64990
AW        -2.38730
AN        18.87860
TX         8.65160
HACA2     -3.09870
CA        61.95360

02:27 Preset to BS9177

02:40 Paul is reverting OSS to the "previous" M3 position for LUCIFER LFBG (what is was on Sep 14-16 UT == Absolute Tip -615, Absolute Tilt -20).

Pointing Check

(numbers didn't make it to the log)

Field Aberration Measurements

/OldRepository/AGW_Data/100921

Problems with Rotator

02:58 LFBG Rotator fault. Steve restarts MCSPU

03:02 Rotator faults again. Steve/Geno go upstairs to press the physical reset button in URTH.

Field Aberration Measurements Resume

03:20 Preset to BS9177

data in /Old/Repository/AGW_Data/100921/field_a

Persson's Star = BS9177
Start time = 03:21

Guide Star wfscimage000*
0 25-28
41 29-31
0 32
57 33-35
0 36-40
32 41-43
0 44
14 45-47
0 48-49
53 50-52
0 53
0 54
51 55-57
0 58
25 59-61
0 62
10 63-65 double!
0 66-67
54 68-70
0 71
46 72-74
0 75-76
15 77-79
0 80-81
16 82-84
0 85-86
0 87-88
41 89-91
0 92
57 93-95
0 96
32 97-99
0 100
14 101-103
0 104-106
0 107-108
53 109-110

End time = 04:55

Note that IDL was sending Zernikes Z4 through Z20 for this dataset, but not Z22.

05:00 Steve moved SX M2 to Tip=-10, Tilt=-5 using PSFGUI

05:10 preset to BS9188

05:12 Preset to ACT0323 to set IE and CA

Adjusting M3 Tip-Tilt

05:15 preset to bs9188 and collimated on-axis.

05:37 preset to bs9188, gs=24 to measure focal plane tilt. Measure ~-300nm Z4 wfsc 13-14

05:40 Steve put +100 Tip M3 in Global Offset then measured ~-600nm Z4 wfsc 15-16

05:45 collimate on-axis wfsc 19-21

05:47 preset to gs=24, Steve remove +100 Tip and put -100 Tip on M3 in Global Offset. Measured -700nm focus wfsc 22-24

05:50 collimate on-axis wfsc 25-33

05:55 measure gs=0, wfsc 34-36 ~-150nm Z4. See about the same on-axis

More Field Aberration Measurements

06:12 Set IE and CA with Lucifer

06:20 Preset to BS9188

data in /OldRepository/AGW_Data/100921/field_b

Persson's Star = BS9188
Start time = 06:15

Guide Star wfscimage000*
0 3-4
41 5-6

See quite a bit of high frequency aberration on primary that could be from our sending Z12-20 earlier. Steve cleared C00 (forces), and that made considerable improvement in the pupils.

Preset on-axis and collimate wfsc 9-11

Guide Star wfscimage000*
0 13-14
34 15-17
0 18-19
31 20-22
0 23
19 24-26
0 27-28
16 29-31
0 32-33
30 34-36
0 37
33 38-40
0 41-42
21 43-45
0 46
32 47-49
0 50
22 51-53
0 54

Moved M3 the wrong direction. Steve moved M3 from the -100 Tip to +75 Tip (+75 Tip in Global Offset corresponds to Absolute -540 Tip)

data in /OldRepository/AGW_Data/100921/field_c

Persson's Star = BS9188
Start time = 06:55

Guide Star wfscimage000*
0 1-3
34 4-6
0 7-8
31 9-11
0 12-13
19 14-16
0 17
16 18-20
0 21
30 22-24
0 25
33 26-28
0 29
21 30-32
0 33-34
32 35-37
0 38

End time = 7:55

Steve added deltas of -30 Tilt and +19 Tip to M2, for total of -35 Tilt and +9 Tip on M2

data in /OldRepository/AGW_Data/100921/field_d

Persson's Star = BS9188
Start time = 08:05

Guide Star wfscimage000*
0 1-2
34 3-5
0 6
31 7-9
0 10
19 11-13
0 14
16 15-17
0 18
30 19-21
0 22-23
33 24-26
0 27-28
21 29-31
0 32-33
32 34-36
0 37
22 38-40
0 41-42

End time = 08:50

09:25 Set d will be our reference set.
M2 Tip=+9 Tilt=-35 arcsec which correspond to IE=-81 CA=+20
M3 Tip=+75 Tilt=0 arcsec (absolute -540 Tip -20 Tilt)

Steve added a delta of +40 Tip to M2 for calibration data (total of +49 Tip)

data in /Old/Repository/AGW_Data/100921/tip_40

Persson's Star = BS9188
Start time = 09:43

Guide Star wfscimage000*
0 13-14
34 15-17
0 18
31 19-21
0 22
19 23-25
0 26
16 27-29
0 30-31
30 32-34
0 35
33 36-38
0 39
21 40-42
0 43
32 44-46
0 47
22 48-50
0 51

End time = 08:50

Steve removed +40 Tip and added +40 Tilt to M2 (total of +5 Tilt) for calibration data

data in /OldRepository/AGW_Data/100921/tilt_40

Persson's Star = BS9107
Start time = 10:45

Guide Star wfscimage000*
0 4
82 5-7
0 8
35 9-11
0 12
19 13-15
0 16
64 17-19
0 20
96 21-23
0 24-25
40 26-28
0 29-30
35 31-33
0 34
48 35-37
0 38
8 39-41
0 42

End time = 11:25

Steve removed +40 Tilt from M2 and added +175 Tilt on M3 for calibration data. Compare this to field_d.

data in /OldRepository/AGW_Data/100921/m3_tilt

Persson's Star = BS9107
Start time = 11:35

Guide Star wfscimage000*
0 1-3
82 4-6
0 7
35 8-10
0 11
19 12-14
0 15
64 16-18
0 19-22
96 23-25
0 26-27
40 28-30
0 31-32
35 33-35
0 36-37
48 38-40
0 41
8 42-44
0 45

End time = 12:30

Range-Balancing Test

12:31 Dave tries his IRAF script to rebalance M1 and M2 ranges with a Mode 1 pointing correction.

To M2, he applied dX=1.787 dY=-0.460 dZ=0 dRX=88.38 dRY=343.70
to get Total Collimation at EL=68 of X=-1.728 Y=-2.503 Z=-1.000 RX=246.08 RY=384.80

To M1, he applied dX=1.636 dY=1.024 dZ=0 DRX=9.05 dRY=55.70
to get Total Collimation of X=0.675 Y=0.894 Z=0.178 RX=53.80 RY=37.085

Close

12:36 Close for impending dawn. T=10.6 degC. Sky is mostly clear, with a few scattered clouds.

-- JohnHill - 21 Sep 2010
Topic attachments
I Attachment Action Size Date Who Comment
daq.20100921.jpgjpg daq.20100921.jpg manage 678 K 22 Sep 2010 - 03:35 JohnHill Plot of Telescope Temperatures
Topic revision: r10 - 22 Sep 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback